OpenStudy (anonymous):

How do i change 1 and 3/8 into a decimal?

OpenStudy (amistre64):

you have to divide

OpenStudy (amistre64):

first change it into an imporoer fraction

OpenStudy (anonymous):

Yes but when you get your remainder what do you do with that?

OpenStudy (amistre64):

\[1\ \frac{3}{8}=\frac{1(8)+3}{8}=\frac{11}{8}\]

OpenStudy (anonymous):

dont stop dividing, just add zeros at the end

OpenStudy (amistre64):

you will end up with no remainder in the end, you just have to push on thru the decimal

OpenStudy (anonymous):

or just use a calculator in dividing :)

OpenStudy (amistre64):

one way to do it, the carpenter way is to reduce your fraction of 1/8 into a decimal 2: .5 4: .25 8: .125 now do 11*.125

OpenStudy (amistre64):

or; 1 + (3*.125)

OpenStudy (anonymous):

1.375 ___________ 8 ) 11.000 8 -- 30 24 -- 60 56 -- 40 40 -- 0

OpenStudy (amistre64):

..... that aint latex :)

OpenStudy (anonymous):

I know! I haven't figured out a good way to do long division with latex yet.

OpenStudy (anonymous):

Oh, but check this out! \[\large (a+b)(c+d) = \underbrace{ac}_\text{First} + \underbrace{bc}_\text{Inner} + \underbrace{ad}_\text{Outer} + \underbrace{bd}_\text{Last}\]

OpenStudy (anonymous):

If i was doing perimeter on a square (it would be the same for everyside). I would just round up the decimal to 1.4?

OpenStudy (anonymous):

haha, i think i meant that the equation would be; 1.4(2) + 1.4(2), right?

OpenStudy (amistre64):

underbrace :)